1

Name  Termin
Ahbay, Cigdem  Fr., 11-13 Uhr, SR 127
Austin, Nathan  Fr., 9-11 Uhr, SR 414
Bargenmeister, Balthasar Fr., 9-11 Uhr, SR 414
Beck, Yvonne  Do., 17-19 Uhr SR 125
Bergner, Johanna  Fr., 9-11 Uhr, SR 125
Biehler, Holger  Fr., 9-11 Uhr, SR 125
Bleichner, Manuel  Fr., 9-11 Uhr, SR 125
Böhmecke, Karoline  Fr., 9-11 Uhr, SR 125
Böttcher, Torsten  Do., 17-19 Uhr SR 125
Breitsch, Milan  Do., 17-19 Uhr SR 125
Bröcheler, Katharina  Do., 17-19 Uhr SR 125
Burbas, Christian  Fr., 11-13 Uhr, SR 127
Büttner, Cynthia  Do., 17-19 Uhr SR 125
Conrad, Tibor  Fr., 11-13 Uhr, SR 127
Eichhorn, Melanie  Do., 17-19 Uhr SR 125
Eipert, Daria  Fr., 9-11 Uhr, SR 125
Eraak, Judith  Fr., 11-13 Uhr, SR 127
Evers, Marie  Fr., 9-11 Uhr, SR 125
Gärtig, John  Fr., 9-11 Uhr, SR 414
Gehring, Melanie  Fr., 9-11 Uhr, SR 125
Gnahm, Senta  Do., 17-19 Uhr SR 125
Goos, Benjamin  Do., 17-19 Uhr SR 125
Haverkampf, Iris  Fr., 11-13 Uhr, SR 127
Härzel, Anne  Fr., 9-11 Uhr, SR 414
Hendel, Peterhans  Fr., 11-13 Uhr, SR 127
Hein, Andreas  Fr., 11-13 Uhr, SR 127
Hiller, Sven  Fr., 9-11 Uhr, SR 414
Imet, Ayari  Fr., 9-11 Uhr, SR 414
Jet Hoe, Tang  Fr., 11-13 Uhr, SR 127
Joos, Martin  Do., 17-19 Uhr SR 125
Krupennyk Halyna, Dolyna Fr., 11-13 Uhr, SR 127
Kurpiers, Andreas  Fr., 11-13 Uhr, SR 127

Kühl, Marlen  Do., 17-19 Uhr SR 125
Lehnerer, Florian  Fr., 9-11 Uhr, SR 414
Marzi, Sarah  Fr., 9-11 Uhr, SR 125
Marquardt, Carla  Do., 17-19 Uhr SR 125
Marquardt, Christian Fr., 9-11 Uhr, SR 414
Matting, Stefan  Fr., 9-11 Uhr, SR 125
Mengersen, Johanna Fr., 11-13 Uhr, SR 127
Metzing, Timo  Fr., 9-11 Uhr, SR 125
Past, Julia  Fr., 11-13 Uhr, SR 127
Rasch, Julia  Fr., 9-11 Uhr, SR 125
Rembold, Philipp  Do., 17-19 Uhr SR 125
Riegger, Julia  Fr., 9-11 Uhr, SR 414
Rosenkranz, Johannes Fr., 9-11 Uhr, SR 414
Sbierski, Jan  Fr., 11-13 Uhr, SR 127
Scharf, Susanne  Fr., 9-11 Uhr, SR 414
Schedensack, Mira  Fr., 9-11 Uhr, SR 125
Schmid, Dorothea  Fr., 9-11 Uhr, SR 414
Schwer, Pascal  Fr., 9-11 Uhr, SR 125
Seither, Daphne  Fr., 9-11 Uhr, SR 414
Stöhlker, Anne-Sophie Fr., 11-13 Uhr, SR 127
Staab, Paul  Do., 17-19 Uhr SR 125
Stelzer, Dominikus  Fr., 11-13 Uhr, SR 127
Stieglitz, Sabine  Do., 17-19 Uhr SR 125
Stirn, Klaus  Fr., 11-13 Uhr, SR 127
Stolz, Jonas  Do., 17-19 Uhr SR 125
Stournaras, Louisa  Fr., 9-11 Uhr, SR 125
Tischer, Katja  Fr., 9-11 Uhr, SR 125
Thoms, Cedric  Fr., 9-11 Uhr, SR 414
Vogelbacher, Georg  Fr., 9-11 Uhr, SR 414
Völkel, Konrad  Fr., 9-11 Uhr, SR 125
Weidner, Jan  Fr., 9-11 Uhr, SR 125
Welzel, Benjamin  Fr., 9-11 Uhr, SR 414
Wolff, Andreas  Do., 17-19 Uhr SR 125